y=3(x+4)^2+1
Vertex, axis of symmetry, x-int, y-int. this graph does not intersect the x=axis
And solve using the quadratic formula

Answer by lwsshak3(11628)

y=3(x+4)^2+1
Vertex, axis of symmetry, x-int, y-int.
**
This is an equation of a parabola that opens upwards.
Its standard form of equation: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of vertex, A=multiplier which affects width or steepness of curve
..
For given equation:y=3(x+4)^2+1
Vertex: (-4,1)
axis of symmetry: x=-4
x-int: none (minimum value=1)
..
y-int:
set x=0, then solve for y
y=3*4^2+1=49
y-int=49